OpenStudy (anonymous):

There are 10 red balls 20 blue balls and 30 green balls in a bag and 5 balls are Selected from it find the probability that all with be blue ? (b) Atleast one will be green

OpenStudy (anonymous):

P(all will be blue) =20C5 / 60C5

OpenStudy (anonymous):

Yes! Good work!

OpenStudy (anonymous):

P(Atleast one green) = 20C1/60C5 * 20C2/60C5 * 20C3/60C5 * 20C4/60C5 * 20C5/60C5 *

OpenStudy (anonymous):

For the second part, instead of adding up P(1) + P(2) + ... + P(5) for at least one green. It is better to get P(no green balls) and subtract that from 1.

OpenStudy (anonymous):

So, 30 "non-greens" in 60.

OpenStudy (anonymous):

So, 1 - 30C5/60C5

OpenStudy (anonymous):

\[1-P(\text{0 are green})=1-(\frac{30}{60}\frac{29}{60}\frac{28}{60}\frac{27}{60}\frac{26}{60}))\]

OpenStudy (anonymous):

Assuming balls are not replaced

OpenStudy (anonymous):

@henpen , that's not right because that formula numerator on the right-side term is going to give you permutations instead of the combinations that you want.

OpenStudy (anonymous):

I think you're overcomplicating.

OpenStudy (anonymous):

Sorry- I made an \[1-(\frac{30}{60}\frac{29}{59}\frac{28}{58}\frac{27}{57}\frac{26}{56})\]error earlier

OpenStudy (anonymous):

Formula is clear and in its most simplified form and given already. You can't get simpler than what I already gave: So, 1 - 30C5/60C5

OpenStudy (anonymous):

I am answering b, by the way

OpenStudy (anonymous):

Yes, so am I, but you are not answering anywhere near correctly.

OpenStudy (anonymous):

Why evoke combinatorics? There is only 1 way for there not to be at least 1 green.

OpenStudy (anonymous):

That is, GGGGG

OpenStudy (anonymous):

\[P(\text{at least 1 green})=1-P(G,G,G,G,G)\]Do you agree with this?

OpenStudy (anonymous):

%15 percent will Blue.

OpenStudy (anonymous):

@henpen ,Your corrected formula will work and can be derived from what I already gave. If there were a denominator of 5! for both numerator and denominator, then they are the same. Now that you corrected your formula, you are on the right track. Same answer. The reason I put the answer in that form is because that is how the soultion is classically represented and is the most straightforward logically-speaking.

OpenStudy (anonymous):

I'm sorry for the error earlier. It just seemed silly to haul in nCr notation where intuition suffices.

OpenStudy (anonymous):

Not, silly. Logical. For those not familiar with the mathematical theory, it is good to reinforce the theory, again, your formula is derivable from mine. It is the same, so no quibbling is necessary here. And the numerical answer for part #2 is then approx. 0.9739

OpenStudy (anonymous):

thxxx guys

OpenStudy (anonymous):

yw!